OpenStudy (anonymous):

2x^2-7x-13=-10 solve graphically show steps

jimthompson5910 (jim_thompson5910):

First add 10 to both sides 2x^2-7x-13=-10 2x^2-7x-13+10=-10+10 2x^2-7x-3=0

jimthompson5910 (jim_thompson5910):

Now graph 2x^2-7x-3 on graph paper or on a graphing calculator. A graphing calculator is much more accurate so I'd use that.

jimthompson5910 (jim_thompson5910):

If you don't have a graphing calculator, you can use something like desmos https://www.desmos.com/calculator

OpenStudy (anonymous):

can i use mathway for the calculature?

jimthompson5910 (jim_thompson5910):

what's the link?

OpenStudy (anonymous):

mathway.com

OpenStudy (anonymous):

i can just use the one u gave me

jimthompson5910 (jim_thompson5910):

I don't see how to graph with mathway, so yeah, use desmos

jimthompson5910 (jim_thompson5910):

graph 2x^2-7x-3 look for the x-intercepts

OpenStudy (anonymous):

0,-3 ?

jimthompson5910 (jim_thompson5910):

that's the y-intercept

jimthompson5910 (jim_thompson5910):

the x-intercepts are the points where the graph crosses the x axis

OpenStudy (anonymous):

oh read it wrong

OpenStudy (anonymous):

-0.39, 0 and 3.89

jimthompson5910 (jim_thompson5910):

I'm getting -0.386 and 3.886

jimthompson5910 (jim_thompson5910):

but it looks like you got roughly the same

OpenStudy (anonymous):

is that the answers or another step

jimthompson5910 (jim_thompson5910):

those are the two solutions

jimthompson5910 (jim_thompson5910):

approximate solutions